# LLMs on the Line: Data Determines Loss-to-Loss Scaling Laws

### v1

@@ Abstract
Scaling laws guide the development of large language models (LLMs) by offering estimates for the optimal balance of model size, tokens, and compute. More recently, loss-to-loss scaling laws that relate losses across pretraining datasets and downstream tasks have emerged as a powerful tool for understanding and improving LLM performance. In this work, we investigate which factors most strongly influence loss-to-loss scaling. Our experiments reveal that the pretraining data and tokenizer determine the scaling trend. In contrast, model size, optimization hyperparameters, and even significant architectural differences, such as between transformer-based models like Llama and state-space models like Mamba, have limited impact. Consequently, practitioners should carefully curate suitable pretraining datasets for optimal downstream performance, while architectures and other settings can be freely optimized for training efficiency.
@@ 1 Introduction
Figure 1: LLMs’ loss-to-loss scaling follows power laws primarily shaped by the choice of pretraining data and tokenizer. Using Llama trained on FineWeb-Edu as a baseline, we intervene on various factors to assess their impact on train-to-test loss scaling. Changing the pretraining data has the largest effect, followed by the choice of tokenizer. Switching the architecture, e.g., from Llama to Mamba, has limited impact, while factors like model size, context length, and optimizer settings exert little-to-no influence.
@@
Brandfonbrener et al. (2024) show that downstream scaling laws can be decomposed into compute-to-train-loss scaling laws and (train)-loss-to-(test)-loss scaling laws. The combination of compute-to-loss and loss-to-loss scaling laws enables efficient and accurate prediction of a model’s downstream performance. Moreover, holistic downstream scaling laws often optimize for a single task or average performance across tasks (Gadre et al., 2024; Schaeffer et al., 2024), whereas loss-to-loss (especially test-to-test) scaling laws can help tune a model’s performance across a broader range of downstream tasks, e.g., to ensure broad or robust generalization.
While the impact of design choices like pretraining distribution, architecture, tokenizer, optimizer settings, etc. on compute-to-loss scaling laws is fairly well understood (Kaplan et al., 2020; Hoffmann et al., 2022; Tay et al., 2022; Wang et al., 2024; Porian et al., 2025; Du et al., 2025), a similar understanding is missing for loss-to-loss scaling laws. To close this gap, we extend the work of Brandfonbrener et al. (2024); Du et al. (2025), which analyze loss-to-loss relationships within a single architectural and training setup. Adding to that, our study systematically explores how multiple factors influence scaling laws across a diverse range of architectures and training configurations.
Our analysis additionally draws inspiration from a body of work in robustness evaluation of vision (and later language) models (Taori et al., 2020; Miller et al., 2021; Fang et al., 2022; Awadalla et al., 2022). These works show that model performance on different distributions is frequently strongly correlated, and most model and training settings have little-to-no impact on the task-to-task scaling trend of model performance. We treat loss-to-loss curves similarly and perform a series of interventions using over 6000 model checkpoints to understand what design choices causally affect scaling trends.
We make three main observations, illustrated in Fig. 1: 1. LLMs’ loss-to-loss scaling consistently follows shifted power laws. 2. Pretraining data and tokenizer are the most salient factors for these scaling laws. 3. In contrast, architecture plays a minor role, while model size, context length, and optimizer settings have negligible impact on loss-to-loss scaling.
Further, we put our observations in the context of downstream scaling laws and discuss the relationship between loss-to-loss and compute-to-loss scaling laws. Our results indicate that different LLM architectures might encode very similar inductive biases, freeing practitioners to optimize architectures for training efficiency without adversely affecting downstream scaling laws.
@@ 2 From Scaling Laws to Interventions
Loss-to-loss scaling laws aim to improve the transferability of scaling insights between training setups by examining the relationship between training (or validation) and test losses, between different validation losses, or between different test losses (Brandfonbrener et al., 2024). This perspective is crucial for several reasons. First, train-to-train (or validation-to-validation) scaling implies how scaling laws transfer across datasets (Brandfonbrener et al., 2024). Second, incorporating train-to-test (or validation-to-test) scaling laws alongside compute-to-train scaling laws provides more precise insight into how compute budgets translate to downstream performance and can help study emergent abilities of models (Du et al., 2025). Third, while compute-to-loss scaling laws often target a single downstream task or average task performance, train-to-test and test-to-test scaling laws can help tune a model’s performance across diverse tasks, e.g., to foster the development of generalist LLMs with a balanced task performance.
@@
Motivated by these insights, we aim to uncover the factors determining loss-to-loss scaling to guide practitioners in developing models for specific downstream performance. Our insights complement the findings from Awadalla et al. (2022), who show that accuracy-accuracy scaling trends in comprehension tasks are agnostic to architecture type (e.g., encoder-only, encoder-decoder, decoder-only) after fine-tuning. In contrast to their study, we focus on zero-shot generalization across a diverse set of tasks, specifically investigating state-of-the-art decoder-only architectures such as GPT (Radford et al., 2019), Llama (Grattafiori et al., 2024), and Mamba (Gu & Dao, 2024; Dao & Gu, 2024).
@@ 3 Fitting Loss-to-Loss Scaling Laws
Figure 2: Loss-to-loss scaling consistently obeys power laws. We extend results from Brandfonbrener et al. (2024) by many architectures, training settings, and validation/test sets. We show illustrative shifted power laws for Mamba trained on FineWeb-Edu here; more configurations and test sets can be found in App. B. For clarity, scatter plots display a random sample of all data points; all points are used to fit the scaling laws.
@@
Brandfonbrener et al. (2024) predict train-to-train and train-to-test scaling laws to follow a shifted power law [^1] Brandfonbrener et al. (2024) do not state a train-to-train scaling law for non-compute-matched data. Our form here follows from their Eq. 4 when assuming an irreducible error as in Eqs. 6, 7.
@@
Given a model configuration, we obtain 200 to 800 by recording checkpoints throughout training, across model sizes, and for various seeds. All points are used to first estimate $E_{x|p},E_{y|p}$ from individual compute-to-loss scaling laws and then fit $K,\kappa$.
Note that to study the impact of various training settings, we show losses of the same model on different datasets on the x- and y-axis. This should not be confused with some curves in Brandfonbrener et al. (2024) that showed the losses of different compute-matched models on the x- and y-axis.
With this setup, we can now analyze the loss-to-loss scaling laws of models trained with different configurations. Brandfonbrener et al. (2024) only showed loss-to-loss scaling a single architecture with a fixed training recipe: Olmo (Groeneveld et al., 2024). We extend their analysis by multiple architectures, pretraining sets, tokenizers, and training settings, all listed in 4. As an illustrative example, we show loss-to-loss scaling for Mamba trained on FineWeb-Edu in Fig. 2; more examples with additional test sets are listed in App. B and throughout 4.
@@
Note also that loss-to-loss scaling follows a power law even for datasets on which the model never reaches high accuracy. E.g., Mamba in Fig. 2 never surpasses chance performance on ARC-Challenge and OpenBookQA, yet Eq. 1 describes the test loss equally well. This underlines the usefulness of loss-to-loss scaling laws to study model behavior.
@@ 4 A Causal Analysis of Loss-to-Loss Scaling
Figure 5: The tokenizer has a moderate impact on loss-to-loss scaling laws. Models are matched on pretraining data and architecture.
@@
Our basic procedure is outlined in Fig. 3. As mentioned in 2, our approach is motivated by similar studies in the robustness literature. In contrast to that setting, we here lack paired in-distribution and out-of-distribution datasets. Instead, we simply consider all combinations of validation and test sets. For ease of visualization when intervening on the pretraining data, we always show FineWeb-Edu validation loss on the x-axis, even for models trained on different pretraining distributions. This choice is arbitrary and does not affect our results; see App. C. Similarly, we here report results for scaling laws of average validation and test loss; results for individual losses can be found in App. D.
@@
We use LM Evaluation Harness framework (Gao et al., 2024) to assess model performance on HellaSwag (Zellers et al., 2019), COPA (Gordon et al., 2011), WinoGrande (Sakaguchi et al., 2019), PIQA (Bisk et al., 2019), OpenBookQA (Mihaylov et al., 2018), SocialIQA (Sap et al., 2019), CommonsenseQA (Talmor et al., 2019), MMLU (Hendrycks et al., 2021), as well as ARC-Easy and ARC-Challenge (Clark et al., 2018).
Architectures
@@ 4.1 Pretraining Data, Tokenizer, and Architecture
Fig. 4 illustrates the substantial impact pretraining data has on loss-to-loss scaling. Across architectures and compute (in different columns), changing the pretraining data leads to a large shift in the loss-to-loss curve. The only exception is the last column, where we compare Hugging Face models trained on The Pile and a deduplicated version. Models trained on either version lie on the same curve, suggesting that the deduplication procedure successfully reduced the dataset size while producing a similar distribution that does not significantly impact loss-to-loss scaling.
@@
Fig. 5 shows that the tokenizer, too, affects loss-to-loss scaling laws, albeit less strongly than pretraining data. It is interesting to see how slight deviations in the tokenizer can have a pronounced effect, particularly for train-to-train scaling laws. While the slight vocabulary size difference between gpt2-HF and gpt-neox has little impact on loss-to-loss scaling (last column), the different handling of special tokens in gpt2 and gpt2-HF does. To the best of our knowledge, this effect has not been observed before and could be explored in future work.
Takeaway 3  With fixed architecture and pretraining data, changing the tokenizer leads to moderate changes in loss-to-loss scaling laws; see Fig. 5.
@@ 4.2 Model Size, Context Length, and Optimization
To provide a more succinct overview, we only show train-to-train scaling laws in this section; additional train-to-test scaling laws for the same intervention can be found in App. E. We also do not show fitted power laws here since we display many more models per plot than in 4.1, and the scaling trends are clearly discernible.
@@
We first examine the influence of model size by training Llama and Mamba models with varying depths and widths (see App. A for details). Fig. 7 shows the results: Despite significant differences in parameter count, the loss-to-loss scaling trends remain unchanged. These findings align well with Du et al. (2025), who observed that model size has little effect on loss-to-loss scaling for GPT models. We extend this conclusion to Llama and Mamba and across multiple pretraining distributions.
@@ 5 Discussion and Future Work
On New Training Paradigms
Our study intentionally focuses on models trained with standard loss functions and conventional training settings to guide practitioners. The limited impact of existing paradigms does not preclude innovative training approaches from improving loss-to-loss scaling. In fact, a recent work by Saunshi et al. (2024) demonstrates that gradually increasing model depth and initializing based on layers from a smaller model produces markedly different scaling behavior, particularly in how perplexity translates to downstream accuracy. Similar structured growth approaches could offer new pathways for improving scaling efficiency and generalization for decoder-only LLMs trained with next-token prediction. We leave this exercise for future work.
@@
Our study clearly distinguishes between factors with substantial and limited impact on loss-to-loss scaling. While our conclusions are inherently shaped by the specific settings we explored, the observed trends provide strong empirical evidence for these distinctions. Given the strong and consistent impact of pretraining data and tokenizer, we can confidently conclude that these interventions affect loss-to-loss scaling. While we observed only a limited impact of the architecture, this effect was also consistent across major state-of-the-art architectures including Llama, GPT, and Mamba — which collectively represent the dominant paradigms in large-scale language modeling. Given this exhaustive set, it is hard to argue that other architectures would meaningfully alter loss-to-loss scaling.
@@
Across the wide range of size configurations (App. A) we test, all models exhibit very consistent loss-to-loss scaling. Similarly, the effect we observed for different context lengths is very consistent within our test range (1024, 2048, 3076), which aligns with commonly used configurations (Black et al., 2021; Wang & Komatsuzaki, 2021; Biderman et al., 2023; Penedo et al., 2024; Black et al., 2022). While we acknowledge the possibility that larger models or longer context lengths could influence loss-to-loss scaling, such an effect — if present — is unlikely. For optimization settings, we again consider configurations widely used in LLM training (Shoeybi et al., 2020; Karpathy, 2022; Videau et al., 2024), including variations in optimizer type, learning rate, weight decay, and scheduling. While our results indicate that these choices do not meaningfully alter loss-to-loss scaling within the explored settings, we acknowledge that the space of optimization techniques is vast, and our list is not exhaustive. It remains possible that a principled optimization strategy, different from current best practices, could induce new scaling behaviors. However, our findings suggest that optimization settings are not a primary driver of loss-to-loss scaling trends within the bounds of conventional language model training.

### v2

@@ Abstract
Scaling laws guide the development of large language models (LLMs) by offering estimates for the optimal balance of model size, tokens, and compute. More recently, loss-to-loss scaling laws that relate losses across pretraining datasets and downstream tasks have emerged as a powerful tool for understanding and improving LLM performance and generalization. In this work, we investigate which factors most strongly influence loss-to-loss scaling. Our experiments reveal that the pretraining data determines the scaling trend. In contrast, model size, optimization hyperparameters, tokenizer and even significant architectural differences, such as between transformer-based models like Llama and state-space models like Mamba, generally have limited impact. Consequently, practitioners should carefully curate pretraining datasets for optimal downstream performance, while architectures and other settings can be freely optimized for training efficiency.
@@ 1 Introduction
Figure 1: LLMs’ loss-to-loss scaling follows power laws primarily shaped by the choice of pretraining data. Using Llama trained on FineWeb-Edu as a baseline, we intervene on various factors to assess their impact on train-to-test loss scaling. Changing the pretraining data has the largest effect. Changing the tokenizer, the architecture (e.g., from Llama to Mamba), model size, context length, and optimizer settings have little-to-no influence. For fair comparison across tokenizers, we normalize loss (negative-log-likelihood, nll) by bytes (bits-per-byte, BPB). We report goodness of fit for the fitted power laws as $R^{2}$. Power laws are fitted using all checkpoints, but only a random subset is shown to avoid visual clutter. Llama-7B checkpoints run for a few steps are highlighted with a black border.
@@
Brandfonbrener et al. (2024) show that downstream scaling laws can be decomposed into compute-to-train-loss scaling laws and (train)-loss-to-(test)-loss scaling laws. The combination of compute-to-loss and loss-to-loss scaling laws enables efficient and accurate prediction of a model’s downstream performance. Moreover, holistic downstream scaling laws often optimize for a single task or average performance across tasks (Gadre et al., 2024; Schaeffer et al., 2024), whereas loss-to-loss (especially test-to-test) scaling laws can help tune a model’s performance across a broader range of downstream tasks, e.g., to ensure broad or robust generalization. Furthermore, train-to-test (or val-to-test) loss scaling laws characterize model generalization behavior. Specifically, these scaling laws are curves depicting the relationship between training or validation loss and downstream test performance, onto which we fit power-law functions.
While the impact of design choices like pretraining distribution, architecture, tokenizer, optimizer settings, etc. on compute-to-loss scaling laws is fairly well understood (Kaplan et al., 2020; Hoffmann et al., 2022; Tay et al., 2022; Wang et al., 2024; Porian et al., 2025; Du et al., 2025), a similar understanding is missing for loss-to-loss scaling laws. To close this gap, we extend the work of Brandfonbrener et al. (2024); Du et al. (2025), which analyze loss-to-loss relationships within a single architectural and training setup. Adding to that, our study systematically explores how multiple factors influence scaling laws across a diverse range of architectures and training configurations. Such an analysis on train-to-test (or val-to-test) scaling laws can help to understand factors that influence model generalization.
Our study draws inspiration from a body of work in robustness evaluation of vision (and later language) models (Taori et al., 2020; Miller et al., 2021; Fang et al., 2022; Awadalla et al., 2022). These works show that model performance on different distributions is frequently strongly correlated, and most model and training settings have little-to-no impact on the task-to-task scaling trend of model performance. We treat loss-to-loss curves similarly and perform a series of interventions using over 6000 model checkpoints to understand what design choices causally affect loss-to-loss scaling laws.
We make three main observations, illustrated in Fig. 1: 1. LLMs’ loss-to-loss scaling consistently follows shifted power laws. 2. Pretraining data is the most salient factor for these scaling laws. 3. In contrast, architecture and tokenizer generally play a minor role, while. model size, context length, and optimizer settings have little-to-no impact on loss-to-loss scaling. These observations imply that if two models with different training setups—but trained on the same data—achieve similar training losses, they will exhibit closely matched downstream test performance.
Our results indicate that common LLM architectures and training setups might encode very similar inductive biases, freeing practitioners to optimize them for training efficiency without adversely affecting downstream generalization.
@@ 2 From Scaling Laws to Interventions
Loss-to-loss scaling laws aim to improve the transferability of scaling insights between training setups by examining the relationship between training (or validation) and test losses, between different validation losses, or between different test losses (Brandfonbrener et al., 2024). Typically, ”training” or ”validation” refers to general-purpose, open-text corpora used for pretraining, whereas ”test” corresponds specifically to downstream tasks. This perspective is crucial for several reasons. First, train-to-train (or validation-to-validation) scaling implies how scaling laws transfer across datasets (Brandfonbrener et al., 2024). Second, incorporating train-to-test (or validation-to-test) scaling laws alongside compute-to-train scaling laws provides more precise insight into how compute budgets translate to downstream performance and can help study emergent abilities of models (Du et al., 2025). Third, while compute-to-loss scaling laws often target a single downstream task or average task performance, train-to-test and test-to-test scaling laws can help tune a model’s performance across diverse tasks, e.g., to foster the development of generalist LLMs with a balanced task performance. Finally—and importantly—train-to-test (or validation-to-test) scaling laws serve as useful tools for understanding generalization.
@@
Motivated by these insights, we aim to uncover the factors determining loss-to-loss scaling to better understand what matters for generalization. Our insights complement the findings from Awadalla et al. (2022), who show that accuracy-accuracy scaling trends in comprehension tasks are agnostic to architecture type (e.g., encoder-only, encoder-decoder, decoder-only) after fine-tuning. In contrast to their study, we focus on zero-shot generalization across a diverse set of tasks, specifically investigating state-of-the-art decoder-only architectures such as GPT (Radford et al., 2019), Llama (Grattafiori et al., 2024), and Mamba (Gu & Dao, 2024; Dao & Gu, 2024).
@@ 3 Fitting Loss-to-Loss Scaling Laws
Figure 2: Loss-to-loss scaling consistently obeys power laws. We extend results from Brandfonbrener et al. (2024) to many architectures, training settings, and validation/test sets. We show illustrative shifted power laws for Mamba trained on FineWeb-Edu here; more configurations and test sets can be found in App. E. For clarity, scatter plots display a random sample of all data points; all points are used to fit the scaling laws.
@@
Brandfonbrener et al. (2024) predict train-to-train and train-to-test scaling laws to follow a shifted power law [^1] Eq. 1 here follows from Brandfonbrener et al. (2024) Eq. (4) when assuming an irreducible error as in Eqs. (6, 7); see App. A.
@@
Given a model configuration, we collect 200 to 800 checkpoints throughout training, across model sizes, and for various seeds. All points are used to first estimate $E_{x|p},E_{y|p}$ from individual compute-to-loss scaling laws and then fit $K,\kappa$. Refer to App. A and App. B for more details.
Note that to study the impact of various training settings, the x- and y-axis show losses on different datasets of the same model. This should not be confused with some figures in Brandfonbrener et al. (2024) where the x- and y-axis show losses of different compute-matched models.
With this setup, we can now analyze the loss-to-loss scaling laws of models trained with different configurations. Brandfonbrener et al. (2024) only showed loss-to-loss scaling a single architecture with a fixed training recipe: Olmo (Groeneveld et al., 2024). We extend their analysis by multiple architectures, pretraining sets, tokenizers, and training settings, all listed in 4. As an illustrative example, we show loss-to-loss scaling for Mamba trained on FineWeb-Edu in Fig. 2; more examples with additional test sets are listed in App. E and throughout 4.
@@
Note also that loss-to-loss scaling follows a power law even for datasets on which the model never reaches high accuracy. E.g., Mamba in Fig. 2 never surpasses chance performance on ARC-Challenge, yet Eq. 1 describes the test loss equally well. This underlines the usefulness of loss-to-loss scaling laws to study model behavior.
@@ 4 A Causal Analysis of Loss-to-Loss Scaling
Figure 5: The tokenizer has a minor impact on loss-to-loss scaling laws. Models are matched on pretraining data and architecture.
@@
Our basic procedure is outlined in Fig. 3. As mentioned in 2, our approach is motivated by similar studies in the robustness literature. In contrast to that setting, here we lack paired in-distribution and out-of-distribution datasets. Instead, we simply consider all combinations of validation and test sets. Typically, ”validation” refers to general-purpose open-text corpora used during pretraining, whereas ”test” corresponds specifically to downstream tasks. For ease of visualization when intervening on the pretraining data, we always show FineWeb-Edu validation loss on the x-axis, even for models trained on different pretraining distributions. This choice is arbitrary and does not affect our results; see App. F. Similarly, we here report results for scaling laws of average validation and test loss; results for individual losses can be found in App. G.
@@
We use LM Evaluation Harness framework (Gao et al., 2024) to assess model performance on HellaSwag (Zellers et al., 2019), COPA (Gordon et al., 2011), WinoGrande (Sakaguchi et al., 2019), PIQA (Bisk et al., 2019), SocialIQA (Sap et al., 2019), CommonsenseQA (Talmor et al., 2019), MMLU (Hendrycks et al., 2021), as well as ARC-Easy and ARC-Challenge (Clark et al., 2018).
Models
@@ 4.1 Pretraining Data, Tokenizer, and Architecture
Fig. 4 illustrates the substantial impact pretraining data has on loss-to-loss scaling. Across architectures and compute (in different columns), changing the pretraining data leads to a large shift in the loss-to-loss curve. Notably, in the last column, we compare Hugging Face models trained on The Pile versus its deduplicated variant. Models trained on these two datasets fall onto slightly different scaling curves, indicating that deduplication has meaningfully altered the underlying distribution and thereby also affecting the loss-to-loss scaling.
@@
Fig. 5 illustrates the impact of tokenizer choice on loss-to-loss scaling laws. To enable fair comparisons between models trained with different tokenizers, we measure performance using bits-per-byte (BPB) Gao et al. (2020). In our experiments, we observe only minor tokenizer-induced variations in scaling curves, as they remain closely aligned. However, the pretrained ’ablation models’ from Penedo et al. (2024) (first and second columns) notably deviate from our curves. One key distinction is that the models from Penedo et al. (2024) utilize ’weight tying’ between the token embeddings in the first layer and the output token predictions in the final layer, whereas our models do not. This difference in training setup could explain the observed discrepancy in the scaling curves. We leave a detailed investigation of this hypothesis for future work.
Takeaway 3  With fixed architecture and pretraining data, changing the tokenizer generally leads to minor changes in loss-to-loss scaling laws; see Fig. 5.
@@ 4.2 Model Size, Context Length, and Optimization
To provide a more succinct overview, we only show train-to-train scaling laws in this section; additional train-to-test scaling laws for the same intervention can be found in App. H. We also do not show fitted power laws here since we display many more models per plot than in 4.1, and the scaling trends are clearly discernible.
@@
We first examine the influence of model size by training Llama and Mamba models with varying depths and widths (see App. C for details). Fig. 7 shows the results: Despite significant differences in parameter count, the loss-to-loss scaling trends remain unchanged. These findings align well with Du et al. (2025), who observed that model size has little effect on loss-to-loss scaling for GPT models. We extend this conclusion to Llama and Mamba and across multiple pretraining distributions.
@@
5 Point-wise Comparisons
In our experiments so far, we have demonstrated that nearly all factors we investigated—except the choice of pretraining data—have minimal to no effect on loss-to-loss scaling. This implies a useful generalization: if two distinct training setups achieve similar training or validation losses on the same pretraining data, they will exhibit similar test losses across many downstream tasks. To explicitly illustrate this point, we present detailed point-wise comparisons between Mamba and Llama models trained on identical datasets in Table 1, highlighting their closely matched downstream performance.
Table 1: Comparison of Mamba and Llama models. Models are compared after seeing approximately 8 Billion Tokens of their pretraining data, at two different model sizes. We observe similar downstream validation and test losses. PT: Pretraining, SPJ: SlimPajama, RW: RefinedWeb, FW-E: FineWeb-Edu, ARC-C: ARC-Challenge, and ARC-E: ARC-Easy.
Model Size PT. Validation Losses Test Task Losses Avg Avg Test
Type Data C4 Pile UC FW-E RW SPJ ARC-C ARC-E HellaSwag MMLU PIQA Val. Loss Task Loss
Mamba 421M FW-Edu 3.66 4.02 2.90 3.77 3.78 5.06 5.27 3.01 2.14 3.94 3.63 3.88
Llama 417M FW-Edu 3.66 3.89 2.90 3.74 3.72 5.26 5.73 3.02 2.24 3.96 3.58 4.04
Mamba 421M Pile UC 3.70 2.75 3.32 3.76 3.50 5.58 6.07 3.23 4.51 4.14 3.41 4.70
Llama 417M Pile UC 3.71 2.75 3.32 3.77 3.49 5.55 5.96 3.26 4.46 4.18 3.41 4.68
Mamba 421M C4 3.17 4.13 3.31 3.85 3.61 5.69 6.34 2.94 5.04 3.73 3.61 4.75
Llama 417M C4 3.17 4.05 3.29 3.82 3.57 5.63 6.25 2.95 5.33 3.76 3.58 4.78
Mamba 368M FW-Edu 3.74 4.15 2.98 3.85 3.86 5.13 5.42 3.10 3.26 4.01 3.72 4.18
Llama 365M FW-Edu 3.74 3.98 2.98 3.84 3.81 5.39 5.78 3.10 3.52 4.04 3.67 4.36
Mamba 368M Pile UC 3.78 2.83 3.40 3.85 3.59 5.75 6.35 3.31 5.08 4.17 3.49 4.93
Llama 365M Pile UC 3.79 2.82 3.41 3.86 3.58 5.97 6.56 3.32 4.96 4.28 3.49 5.02
Mamba 368M C4 3.25 4.23 3.39 3.92 3.69 5.92 6.66 3.03 5.45 3.79 3.70 4.97
Llama 365M C4 3.27 4.13 3.38 3.91 3.65 5.77 6.33 3.05 5.56 3.90 3.67 4.92
@@ 6 Discussion and Future Work
On Other Training Paradigms
Our study intentionally focuses on models trained with standard loss functions and conventional training settings to guide practitioners. The limited impact of existing paradigms does not preclude innovative training approaches from improving loss-to-loss scaling. In fact, a recent work by Saunshi et al. (2024) demonstrates that gradually increasing model depth and initializing based on layers from a smaller model produces markedly different scaling behavior, particularly in how perplexity translates to downstream accuracy. Similar structured growth approaches could offer new pathways for improving scaling efficiency and generalization for decoder-only LLMs trained with next-token prediction. Additionally, in 4.1 we observed that weight tying could play a potential role in shifting these generalization curves. We leave these analyses for future work.
@@
Our study clearly distinguishes between factors with substantial and limited impact on loss-to-loss scaling. While our conclusions are inherently shaped by the specific settings we explored, the observed trends provide strong empirical evidence for these distinctions. Given the strong and consistent impact of pretraining data, we can confidently conclude that this intervention affects loss-to-loss scaling. While we observed only a limited impact of the architecture, this effect was also consistent across major state-of-the-art architectures including Llama, GPT, and Mamba — which collectively represent the dominant paradigms in large-scale language modeling. Given this exhaustive set, it is hard to argue that other architectures would meaningfully alter loss-to-loss scaling.
@@
Across the wide range of size configurations (App. C) we test, all models exhibit very consistent loss-to-loss scaling. Similarly, the effect we observed for different context lengths is very consistent within our test range (1024, 2048, 3076), which aligns with commonly used configurations (Black et al., 2021; Wang & Komatsuzaki, 2021; Biderman et al., 2023; Penedo et al., 2024; Black et al., 2022). While we acknowledge the possibility that larger models or longer context lengths could influence loss-to-loss scaling, such an effect — if present — is unlikely. For optimization settings, we again consider configurations widely used in LLM training (Shoeybi et al., 2020; Karpathy, 2022; Videau et al., 2024), including variations in optimizer type, learning rate, weight decay, and scheduling. While our results indicate that these choices do not meaningfully alter loss-to-loss scaling within the explored settings, we acknowledge that the space of optimization techniques is vast, and our list is not exhaustive. It remains possible that a principled optimization strategy, different from current best practices, could induce new scaling behaviors. However, our findings suggest that optimization settings are not a primary driver of loss-to-loss scaling trends within the bounds of conventional language model training.